OpenStudy (anonymous):

How many solutions does the following equation have? log(3x − 3) = log(x + 1) + log 4

hartnn (hartnn):

on the right side, use the log property that \(\log A +\log B = \log AB\)

OpenStudy (ranga):

Make use of the following:\[\Large \log(A) + \log(B) = \log(AB)\]\[\Large If \quad \log(C) = \log(D) \quad then \quad C = D\]

OpenStudy (anonymous):

this looks like a lot of letters… can you dumb it down? :)

OpenStudy (ranga):

The first one for example says: log(5) + log(6) = log(5*6) = log(30)

hartnn (hartnn):